OpenStudy (sloppycanada):

Write an appropriate inverse variation equation if y = 2 when x = 9

OpenStudy (loser66):

hey, why are you stuck? what is the inverse standard form?

OpenStudy (sloppycanada):

Well 1/x?

OpenStudy (loser66):

not, 1/x, it is k/x

OpenStudy (loser66):

y = k/x = 2 and x =9, then k =?

OpenStudy (sloppycanada):

18

OpenStudy (loser66):

yup, dat sit!! y = 18/x